OpenStudy (anonymous):

[2.07] Solve 3|2x − 1| − 4 = 29 for x. Show your work.

OpenStudy (anonymous):

@dpaInc Thamks For Viewing My Question.

OpenStudy (anonymous):

np... do you know what to do first here?...

OpenStudy (anonymous):

I'm Not Sure.

OpenStudy (anonymous):

let's isolate the absolute value by adding 4 then dividing by 3 to both sides of the equation...

OpenStudy (anonymous):

Ok

OpenStudy (anonymous):

\(\large 3|2x-1|-4=29 \) \(\large 3|2x-1|-4\color {red}{+4}=29\color {red}{+4} \) \(\large 3|2x-1|=33 \) \(\huge \frac{3|2x-1|}{\color {red} {3}}=\frac{33}{\color {red}{3}} \) \(\large |2x-1|=11 \)

OpenStudy (anonymous):

now that we've isolated the absolute value on one side, we'll need to solve these two equations: \(\large 2x-1=11 \) OR \(\large 2x-1=-11 \) can you take it from here?

OpenStudy (anonymous):

@Deldrickg are you in FLVS?